Every day the Operations Team in the UK gets sent a report on the day’s actitivies compiled by Brian Newham, including weather conditions, distance travelled and current position.

In the spirit of sharing, here’s today’s:

Best wishes,

Hugh – Operations HQ

Daily Report

Sunday 10th Feb 2013

Current position S70 33′ 36.7″” E023 28′ 05.1″ (same as yesterday)

Distance travelled today zero km

Weather at 1700 GMT
Wind – SE 25 kts
Temp – minus 3C
Cloud – obscured
Current weather – drifting snow
Visibility – 200m
Contrast – Nil

Today’s action
It was blowing 40kts this morning and that has gradually eased to 25kts. Poor visibility and whiteout conditions. No movement for the Ice Train.

Intentions for tomorrow
There is moderate drift accumulation around camp which will require dozing before we can link units back together and getting moving again. The current improving trend in the weather suggests that we may be able to move tomorrow.

Long term intentions
If all is well we will continue south.